When three coins are tossed,the probability distribution for the random variable x representing the number of heads that occurs is given below compute the variance and standard deviation of the probability distribution


1
Expert's answer
2022-05-16T15:40:20-0400
"\def\arraystretch{1.5}
 \begin{array}{c:c}
 & x \\ \hline
 TTT & 0 \\
 \hdashline
 TTH & 1\\
 \hdashline
 THT & 1 \\
 \hdashline
 HTT & 1 \\
 \hdashline
 THH & 2 \\
 \hdashline
 HTH & 2\\
 \hdashline
 HHT & 2\\
 \hdashline
 HHH & 3 \\
 \hdashline
\end{array}"


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & 0 & 1 & 2 & 3 \\ \hline
 p(x) & 1/8 & 3/8 & 3/8 & 1/8 \\
\end{array}"

"E(X)=\dfrac{1}{8}(0)+\dfrac{3}{8}(1)+\dfrac{3}{8}(2)+\dfrac{1}{8}(3)=\dfrac{3}{2}"

"E(X^2)=\dfrac{1}{8}(0)^2+\dfrac{3}{8}(1)^2+\dfrac{3}{8}(2)^2+\dfrac{1}{8}(3)^2=3"

"Var(X)=\sigma^2=E(X^2)-(E(X))^2"

"=3-(\dfrac{3}{2})^2=\dfrac{3}{4}"

"\sigma=\sqrt{\sigma^2}=\sqrt{\dfrac{3}{4}}=\dfrac{\sqrt{3}}{2}\approx0.8660"